OpenStudy (anonymous):

What is the value of the expression –7 – 5 + 4? A. –12 B. –16 C. –6 D. –8

OpenStudy (unanimoose):

What is -7 - 5?

OpenStudy (anonymous):

-7-5+4

OpenStudy (unanimoose):

My question was: solve -7 - 5 first.

OpenStudy (anonymous):

D) -8 -7-5=-12 -12+4 = -8

OpenStudy (unanimoose):

And then add 4 to your solution.

OpenStudy (unanimoose):

Or just get a straight answer, that works too.

OpenStudy (anonymous):

what is the answer

OpenStudy (unanimoose):

>___<

OpenStudy (anonymous):

ill give medal to who tells me

OpenStudy (anonymous):

look at my post i gave you the answer and how i got to it

OpenStudy (unanimoose):

She just said it.

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

sorry my badsssss